5. Find the absolute maximum and absolute minimum values of the function f(x) = x.ea) on the interval -2 33 32.
Answer #1

plug in all critical numbers and end points of interval in of (X). Highest value among them results absolute marina I leart veto 1 1+2=0 =121= undefined ~=? 22 is critical number. Now plug in x = -222, 22, 2 in of(=mes *-=f(2)=3 = - 3F05518) x==2= f- 3 ze is leart value occurs at x=1-3 results absolute minora The 15 is highest value occurs at n=120 results absolute masina